# Ultra-Slow-Roll Inflation on the Lattice: Backreaction and Nonlinear Effects

### v2

@@ II.3.2 Action for perturbations with and without Wands duality
Figure 4: Top panel: Evolution of $\nu^{2}$ as a function of number of $e$-folds across SR-USR-SR transitions in the three cases considered in this work, with $\Delta N=2$ in each case. While all cases presents a sizeable evolution of $\nu^{2}$ at the onset of USR, in case I, $\nu^{2}$ is nearly constant from deep in the USR phase. Bottom panel: first derivative of $\nu^{2}$ with respect to number of efolds. By construction, at the end of USR around $N=2$, it is ${\cal O}(1)$ and positive (resp. negative) in case II (resp. case III), which makes the former a “repulsive” scenario while the latter can be consider as “attractive”. See the main text for an explanation of the terminology adopted here. By contrast, in case I with approximate Wands duality, $(\nu^{2})^{\prime}$ is two orders of magnitude smaller than the other cases at $N=2$, corresponding to an approximately free theory.
@@
We realise this scenario, albeit approximately, in case I. We consequently fix $\eta$ in phase II and III, while considering a symmetric transition between them such that it retains a nearly constant $\nu$ after the onset of USR, and crucially during the transition between USR and the subsequent SR. Therefore, we only expect a backreaction on the background evolution and no effect on the spectrum of $\delta\phi$ perturbations in case I. Notice that this property is independent from the assumed duration of USR.
We show the evolution of $\nu^{2}$ as a function of number of $e$-folds in Fig. 4. We can notice that in all cases, $\nu^{2}$ transitions from the value $\approx 9/4$ to negative during the onset of USR, while relaxes back to $\approx 4$ afterwards. The final value is obtained by inserting $\eta_{\rm III}=-0.5$ into Eq. (12). Case I approaches a nearly constant value already during USR, suppressing interactions during the last stages, including at the USR-SR transition. In the other cases, the time evolution remains sizeable, with $\nu^{2}$ increasing in case II, and decreasing in case III, at the critical period corresponding to the end of the USR phase (around $N\sim 2$). This corresponds to, respectively, a positive and negative cubic coupling $V_{3}$ in (11b), conventionally referred to as repulsive and attractive interactions. Therefore, in the following, we will refer to case II as the repulsive case and case III as the attractive case. We will comment more precisely on the concrete meaning of these names when analyzing snapshots of lattice simulations of the three cases in Fig. 8.
@@ III.2 Simulation results
In this section we show simulation results for the different potentials introduced in Sec. II, namely case I, II and III of Table 1. For each of these cases, we run several lattice simulations for different potentials that are predicted to generate tree-level power spectra with a peak in the range $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-4}\div 1$. In all plots presented in the following, each line from blue to yellow corresponds to a unit jump of $\log_{10}(\mathcal{P}^{\rm max}_{\zeta,\rm tree})$.
@@ III.2.2 Fluctuations
In the middle panels of Figs. 5, 6, and 7, we show the power spectra of $\zeta$ and $\phi$, and we compare them with the linear theory predictions. For $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-4}$, both $\mathcal{P}_{\phi}$ and $\mathcal{P}_{\zeta}$ match linear predictions in all three cases. For larger $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, however, corrections appear, with O(1) deviations in $\mathcal{P}_{\zeta}$ for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}\geq 0.1$, and deviations exceeding $\sim 10\%$ for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-2}$.555 One should be aware that the spectrum of perturbations at low $k$ in Figs. 5, 6, and 7 are largely affected by sample variance due to the limited number of modes contained in the simulation box.
@@
  • In case I, structures in the inflaton field remain unchanged, confirming the nearly free nature due to the approximate Wands duality.
@@
Eventually, it is also interesting to look at the 1-point probability density function (PDF) of the inflaton fluctuation at the final simulation time, shown in Fig. 9 for all three cases. Contrary to the snapshots, this does not carry any morphological information, but it gives a straightforward visualization of some statistical properties. Like in Fig. 8, we plot the distribution with the $x-$axis normalised to the perturbation variance $\sqrt{\langle\delta\phi^{2}\rangle}$. Therefore, any deviation from the blue curve, for which nonlinear effects are minimal and negligible, indicate non-Gaussianity of the distribution. In case I, the statistics remain nearly Gaussian for all values of $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, consistently with the nearly-free nature of this model at the relevant times of the dynamics. In cases II and III, non-Gaussianity grows with increasing $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$. In case II, this manifests as an increasingly negatively skewed distribution, a suppression of the right tail with $\delta\phi>0$, and a (more modest) enhancement of the left tail with $\delta\phi<0$. This behavior of the tails qualitatively agrees with our discussion on the repulsive interaction, but we stress that the latter aspect is better identified with the morphological information visible in the snapshots, and that the tails, by definition, are also sensitive to higher-order couplings. Similar remarks hold for case III with inverse trends: the larger $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, the more positively skewed the distribution, the smaller the left-tail and the larger the right tail. We highlight that the striking behavior of the right-tail in the orange most extreme scenario (with $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=1$) is a manifestation of a phenomenon that is qualitatively different than the other situations. This tail reflects that, while most patches manage to escape the USR phase, some lack sufficient kinetic energy to climb the potential maximum and instead roll back toward the minimum. This is illustrated by the animation of the simulations (link), where we show the inflaton 1-point PDF moving along the potential. These patches, trapped in a false vacuum, will form PBH after inflation, as described in Caravano et al. (2024); Inomata et al. (2022); Atal et al. (2019, 2020); Escrivà et al. (2023); Flores et al. (2024). By contrast, when $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=0.1$, the exponential behavior of the tail does not correspond to the trapping phenomenon (with the exception of a few extreme points that are not statistically significant).
@@ III.2.3 Discussion
At this stage, two remarks are important. First, the effects of gradients discussed above are not taken into account in the separate universe picture. Second, if this interpretation is correct, the additional acceleration of the inflaton trajectory is temporary and ceases once most points have crossed the USR phase (consistently with Eq. (24) once the $V^{\prime\prime}$ and $V^{(3)}$ terms have become negligible). Thus, backreaction should manifest as an advancement in the background motion, but the background trajectory $\dot{\phi}(\phi)$ should fall back to the same slow-roll attractor in the second SR phase, as already suggested by results in the top right plots of Figs. 5-7. This is confirmed by the simulations, as we present in Fig. 10, where we show that the phase space trajectory $\dot{\phi}(\phi)$ coincides with the homogeneous prediction of SPT without backreaction sufficiently before and after the USR phase, in all cases considered.
Finally, let us discuss the backreaction on $\dot{\phi}$ in relation to the evolution of different energy contributions during the USR phase. In Fig. 11, we show the evolution of the different relative contributions to the energy density in the lattice simulation for the three different cases. When linear perturbation theory holds (e.g., $\mathcal{P}_{\zeta}=10^{-4}$), there is a clear hierarchy among the energy contributions, $V_{\phi}\gg K_{\phi}\gg G_{\phi}$, similar to standard slow-roll inflation. Backreaction becomes significant when the kinetic energy decreases to the level of the gradient energy, $K_{\phi}\sim G_{\phi}$, which is not taken into account in SPT. [^6] Note that $K_{\phi}\sim G_{\phi}$ also at the starting time $N=-2$, when the simulation box is sub-Hubble. This is common in lattice simulations where one initializes sub-Hubble quantum fluctuations as stochastic noise. Contrarily to what happens at $N\simeq 2$, this does not affect the dynamics of the field, which on sub-Hubble scales is trivial and is dominated by the gradient term in the equation of motion. Therefore, at the initial time, the interplay between gradients and other energy contributions is irrelevant. This is confirmed by the agreement between the lattice simulation and linear theory at early times. Moreover, we have seen that the kinetic energy is larger than the one predicted by SPT, suggesting that backreaction effectively protects the kinetic energy from dropping below the gradient energy density.
@@ IV Conclusions and outlook
In this work, we have studied the nonlinear dynamics and backreaction effects in inflation scenarios featuring a transient USR phase using lattice simulations. We showed that, already for power spectra peaking at $\mathcal{P}_{\zeta}^{\rm max}\sim 10^{-2}$, nonlinear effects lead to corrections of the order of 10% on the power spectral amplitude, in agreement with naive order of magnitude expectation of perturbation theory. Importantly though, these corrections occur without violating the conditions for perturbativity, suggesting analytical perturbative computations beyond tree level could be sufficient to describe these models, once the important backreaction effect pointed out in this work is taken into account. As USR inflationary models are particularly relevant for scenarios predicting the formation of PBHs and SIGWs, controlling these nonlinear corrections is crucial for making robust predictions.
@@
We have also shown that in scenarios characterized by an approximate Wands duality, nonlinear self-interactions remain suppressed, and the system behaves nearly as a free theory. In these cases, while backreaction effects on the background evolution are still present, the spectrum of scalar perturbations remains unaffected by nonlinear corrections, as previously suggested in the literature. However, for models that break Wands duality, we observed more pronounced nonlinear interactions, leading to significant non-Gaussian features in the statistics of the inflaton perturbations. In particular, we found that for repulsive cases (characterized by a positive cubic interaction), the distribution of field fluctuations exhibits a suppression of large positive fluctuations, while for attractive cases (with negative cubic interactions), the field develops a heavy exponential tail. This behavior is critical for understanding the potential formation of PBHs, as the tail of the distribution directly affects the probability of generating large overdensities that could collapse into black holes.
@@
It was recently suggested that USR dynamics advocated by some of the PBH formation scenarios may violate perturbativity and induce loop corrections affecting much longer modes associated to the scales observed through the CMB Kristiano and Yokoyama (2024c). [^7] Instead, see Fumagalli et al. (2024) for effects at near infrared scales when the power spectrum enhancement is due to particle production. While the existence and magnitude of this effect is still under debate Cheng et al. (2022); Riotto (2023a); Kristiano and Yokoyama (2024d); Riotto (2023b); Kristiano and Yokoyama (2024b); Franciolini et al. (2024b); Davies et al. (2024); Firouzjahi and Riotto (2024); Iacconi et al. (2024); Motohashi and Tada (2023); Tasinato (2024, 2023); Firouzjahi (2023); Fumagalli (2023); Cheng et al. (2024); Tada et al. (2024); Firouzjahi (2024); Braglia and Pinol (2024); Kawaguchi et al. (2024a); Ballesteros and Egea (2024); Inomata (2024); Fumagalli (2024); Kawaguchi et al. (2024b); Green and Gupta (2024), we are not able to address this question due to the relatively small range of scales that can be tracked in the lattice box. However, we are in the position to comment on some of the related issues. In particular, our result shows perturbativity is still not completely violated as long as the amplitude of perturbations are smaller than order unity (contrary to what happens in oscillatory potentials, where perturbativity is completely broken already for $\mathcal{P}_{\zeta,\rm tree}^{\rm max}=10^{-2}$ Caravano et al. (2024)). Also, we confirm the existence of important backreaction on the background dynamics, typically captured by tadpole contributions, which were suggested to play a crucial role in protecting long modes from loop corrections Inomata (2024); Fumagalli (2024).
@@
Looking forward, it would be interesting to extend the present work along different directions. Firstly, it will be important to derive the full non-Gaussian statistics for the curvature perturbation, including the nonlinear relation between $\zeta$ and $\delta\phi$ from our lattice simulations, to assess whether positive NGs in $\zeta$ are obtained in all scenarios considered here, as suggested in Ref. Firouzjahi and Riotto (2023). Other groups attempted to study the effect of backreaction and nonlinearities in the USR scenarios employing full in-in computations, $\delta N$, or the stochastic inflation formalism (see e.g. Pattison et al. (2017); Biagetti et al. (2018); Ezquiaga et al. (2020); Ballesteros et al. (2020); Pattison et al. (2021); Cai et al. (2022); Biagetti et al. (2021); Figueroa et al. (2022); Raatikainen et al. (2024); Mizuguchi et al. (2024); Ballesteros and Egea (2024); Animali and Vennin (2024); Jackson et al. (2024)), and it would be interesting to compare these techniques with lattice results. Additionally, evaluating the four-point function of the field fluctuations $\delta\phi$ would provide key insights into the non-perturbative source of SIGW (complementing recent efforts in this direction Cai et al. (2019); Unal (2019); Yuan and Huang (2021); Atal and Domènech (2021); Adshead et al. (2021); Abe et al. (2023); Chang et al. (2023); Garcia-Saenz et al. (2023); Li et al. (2023); Perna et al. (2024); Ruiz and Rey (2024)). These next steps will allow us to better quantify the importance of nonlinearities in the USR scenarios, paving the way for robust observational predictions for GW experiments, as well as constraints on the PBH abundance.

### v3

@@ II.3.2 Action for perturbations with and without Wands duality
Figure 4: Top panel: Evolution of $\nu^{2}$ as a function of number of $e$-folds across SR-USR-SR transitions in the three cases considered in this work, with $\Delta N=2$ in each case. While all cases presents a sizeable evolution of $\nu^{2}$ at the onset of USR, in case I, $\nu^{2}$ is nearly constant from deep in the USR phase. Bottom panel: first derivative of $\nu^{2}$ with respect to number of efolds. By construction, at the end of USR around $N=2$, it is ${\cal O}(1)$ and positive (resp. negative) in case II (resp. case III), which makes the former a “repulsive” scenario while the latter can be considered as “attractive”. See the main text for an explanation of the terminology adopted here. By contrast, in case I with approximate Wands duality, $(\nu^{2})^{\prime}$ is two orders of magnitude smaller than the other cases at $N=2$, corresponding to an approximately free theory.
@@
We realise this scenario, albeit approximately, in case I. We consequently fix $\eta$ in phases II and III, while considering a symmetric transition between them such that it retains a nearly constant $\nu$ after the onset of USR, and crucially during the transition between USR and the subsequent SR. Therefore, we only expect a backreaction on the background evolution and no effect on the spectrum of $\delta\phi$ perturbations in case I. Notice that this property is independent from the assumed duration of USR.
We show the evolution of $\nu^{2}$ as a function of number of $e$-folds in Fig. 4. We can notice that in all cases, $\nu^{2}$ transitions from the value $\approx 9/4$ to negative during the onset of USR, while relaxes back to $\approx 4$ afterward. The final value is obtained by inserting $\eta_{\rm III}=-0.5$ into Eq. (12). Case I approaches a nearly constant value already during USR, suppressing interactions during the last stages, including at the USR-SR transition. In the other cases, the time evolution remains sizeable, with $\nu^{2}$ increasing in case II, and decreasing in case III, at the critical period corresponding to the end of the USR phase (around $N\sim 2$). This corresponds to, respectively, a positive and negative cubic coupling $V_{3}$ in (11b), conventionally referred to as repulsive and attractive interactions. Therefore, in the following, we will refer to case II as the repulsive case and case III as the attractive case. We will comment more precisely on the concrete meaning of these names when analyzing snapshots of lattice simulations of the three cases in Fig. 8.
@@ III.2 Simulation results
In this section we show simulation results for the different potentials introduced in Sec. II, namely cases I, II and III of Table 1. For each of these cases, we run several lattice simulations for different potentials that are predicted to generate tree-level power spectra with a peak in the range $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-4}\div 1$. In all plots presented in the following, each line from blue to yellow corresponds to a unit jump of $\log_{10}(\mathcal{P}^{\rm max}_{\zeta,\rm tree})$.
@@ III.2.2 Fluctuations
In the middle panels of Figs. 5, 6, and 7, we show the power spectra of $\zeta$ and $\phi$, and we compare them with the linear theory predictions. For $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-4}$, both $\mathcal{P}_{\phi}$ and $\mathcal{P}_{\zeta}$ match linear predictions in all three cases. For larger $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, however, corrections appear, with O(1) deviations in $\mathcal{P}_{\zeta}$ for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}\geq 0.1$, and deviations exceeding $\sim 20\%$ for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=10^{-2}$.555 One should be aware that the spectrum of perturbations at low $k$ in Figs. 5, 6, and 7 are largely affected by sample variance due to the limited number of modes contained in the simulation box.
@@
  • In case I, structures in the inflaton field remain unchanged, confirming the nearly-free nature due to the approximate Wands duality.
@@
Eventually, it is also interesting to look at the 1-point probability density function (PDF) of the inflaton fluctuation at the final simulation time, shown in Fig. 9 for all three cases. Contrary to the snapshots, this does not carry any morphological information, but it gives a straightforward visualization of some statistical properties. Like in Fig. 8, we plot the distribution with the $x-$axis normalised to the perturbation variance $\sqrt{\langle\delta\phi^{2}\rangle}$. Therefore, any deviation from the blue curve, for which nonlinear effects are minimal and negligible, indicate non-Gaussianity of the distribution. In case I, the statistics remain nearly Gaussian for all values of $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, consistent with the nearly-free nature of this model at the relevant times of the dynamics. In cases II and III, non-Gaussianity grows with increasing $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$. In case II, this manifests as an increasingly negatively skewed distribution, a suppression of the right tail with $\delta\phi>0$, and a (more modest) enhancement of the left tail with $\delta\phi<0$. This behavior of the tails qualitatively agrees with our discussion on the repulsive interaction, but we stress that the latter aspect is better identified with the morphological information visible in the snapshots, and that the tails, by definition, are also sensitive to higher-order couplings. Similar remarks hold for case III with inverse trends: the larger $\mathcal{P}^{\rm max}_{\zeta,\rm tree}$, the more positively skewed the distribution, the smaller the left-tail and the larger the right tail. We highlight that the striking behavior of the right-tail in the orange most extreme scenario (with $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=1$) is a manifestation of a phenomenon that is qualitatively different than the other situations. This tail reflects that, while most patches manage to escape the USR phase, some lack sufficient kinetic energy to climb the potential maximum and instead roll back toward the minimum. This is illustrated by the animation of the simulations (link), where we show the inflaton 1-point PDF moving along the potential. These patches, trapped in a false vacuum, will form PBH after inflation, as described in Caravano et al. (2024); Inomata et al. (2022); Atal et al. (2019, 2020); Escrivà et al. (2023); Flores et al. (2024). By contrast, for $\mathcal{P}^{\rm max}_{\zeta,\rm tree}=0.1$, the exponential tail reflects intrinsic non-Gaussianity of the inflaton field rather than the trapping phenomenon, except for a few statistically insignificant extreme points.
@@ III.2.3 Discussion
At this stage, two remarks are important. First, the effects of gradients discussed above are not taken into account in the separate universe picture. Second, if this interpretation is correct, the additional acceleration of the inflaton trajectory is temporary and ceases once most points have crossed the USR phase (consistent with Eq. (24) once the $V^{\prime\prime}$ and $V^{(3)}$ terms have become negligible). Thus, backreaction should manifest as an advancement in the background motion, but the background trajectory $\dot{\phi}(\phi)$ should fall back to the same slow-roll attractor in the second SR phase, as already suggested by results in the top right plots of Figs. 5-7. This is confirmed by the simulations, as we present in Fig. 10, where we show that the phase space trajectory $\dot{\phi}(\phi)$ coincides with the homogeneous prediction of SPT without backreaction sufficiently before and after the USR phase, in all cases considered.
Finally, let us discuss the backreaction on $\dot{\phi}$ in relation to the evolution of different energy contributions during the USR phase. In Fig. 11, we show the evolution of the different relative contributions to the energy density in the lattice simulation for the three different cases. When linear perturbation theory holds (e.g., $\mathcal{P}_{\zeta}=10^{-4}$), there is a clear hierarchy among the energy contributions, $V_{\phi}\gg K_{\phi}\gg G_{\phi}$, similar to standard slow-roll inflation. Backreaction becomes significant when the kinetic energy decreases to the level of the gradient energy, $K_{\phi}\sim G_{\phi}$, which is not taken into account in SPT. [^6] Note that $K_{\phi}\sim G_{\phi}$ also at the starting time $N=-2$, when the simulation box is sub-Hubble. This is common in lattice simulations where one initializes sub-Hubble quantum fluctuations as stochastic noise. Contrary to what happens at $N\simeq 2$, this does not affect the dynamics of the field, which on sub-Hubble scales is trivial and is dominated by the gradient term in the equation of motion. Therefore, at the initial time, the interplay between gradients and other energy contributions is irrelevant. This is confirmed by the agreement between the lattice simulation and linear theory at early times. Moreover, we have seen that the kinetic energy is larger than the one predicted by SPT, suggesting that backreaction effectively protects the kinetic energy from dropping below the gradient energy density.
@@ IV Conclusions and outlook
In this work, we studied the nonlinear dynamics and backreaction effects in inflationary scenarios featuring a transient USR phase using lattice simulations. We demonstrated that for power spectra peaking at $\mathcal{P}_{\zeta}^{\rm max}\sim 10^{-2}$, nonlinear effects introduce $\sim 20$% corrections to the power spectral amplitude, consistent with the order-of-magnitude expectation from perturbation theory of $O(10\%)$. Accurately accounting for these corrections is essential to make reliable predictions for the generation of PBHs and SIGWs. Interestingly, the size of these effects suggests that perturbativity is not completely violated (in contrast to what happens, for example, in Caravano et al. (2024)). This indicates that perturbative computations beyond tree level—when incorporating the key backreaction effects identified here—could provide a viable analytical description. These effects would appear as large loop corrections around the SPT background, which should be resummed into a backreacted background around which perturbation theory should be performed.
@@
We have also shown that in scenarios characterized by an approximate Wands duality, nonlinear self-interactions remain suppressed, and the system behaves nearly as a free theory. In these cases, while backreaction effects on the background evolution are still present, the spectrum of scalar perturbations remains unaffected by nonlinear corrections, as previously suggested in the literature. However, for models that break Wands duality, we observed more pronounced nonlinear interactions, leading to significant non-Gaussian features in the statistics of the inflaton perturbations. In particular, we found that for repulsive cases (characterized by a positive cubic interaction), the distribution of field fluctuations exhibits a suppression of large positive fluctuations, while for attractive cases (with negative cubic interactions), the field develops a heavy exponential tail. This behavior is critical for understanding the potential formation of PBHs, as the tail of the distribution directly affects the probability of generating large overdensities that could collapse into black holes. Crucially, the exponential tail observed in this work reflects the intrinsic non-Gaussianity of the inflaton field and not that of $\zeta$. Therefore, this tail cannot be directly compared to those arising from the nonlinear relation between $\delta\phi$ and $\zeta$ or from stochastic effects (see, e.g., Cai et al. (2018); Ezquiaga et al. (2020)). Note that, in contrast to our method, these approaches neglect the intrinsic non-Gaussianity prior to the crossing of the coarse-graining scale.
@@
It was recently suggested that USR dynamics advocated by some of the PBH formation scenarios may violate perturbativity and induce loop corrections affecting much longer modes associated with the scales observed through the CMB Kristiano and Yokoyama (2024c). [^7] Instead, see Fumagalli et al. (2024) for effects at near infrared scales when the power spectrum enhancement is due to particle production. While the existence and magnitude of this effect are still under debate Cheng et al. (2022); Riotto (2023a); Kristiano and Yokoyama (2024d); Riotto (2023b); Kristiano and Yokoyama (2024b); Franciolini et al. (2024b); Davies et al. (2024); Firouzjahi and Riotto (2024); Iacconi et al. (2024); Motohashi and Tada (2023); Tasinato (2024, 2023); Firouzjahi (2023); Fumagalli (2023); Cheng et al. (2024); Tada et al. (2024); Firouzjahi (2024); Braglia and Pinol (2024); Kawaguchi et al. (2024a); Ballesteros and Egea (2024); Inomata (2024); Fumagalli (2024); Kawaguchi et al. (2024b); Green and Gupta (2024), we can not directly address this question due to the relatively small range of scales that can be tracked in the lattice box. However, we demonstrated the existence of significant backreaction on the background dynamics, typically captured by tadpole contributions, which have been suggested to play a crucial role in protecting long modes from loop corrections Inomata (2024); Fumagalli (2024).
@@
Looking forward, it would be interesting to extend the present work in different directions. First, it will be important to derive the full non-Gaussian statistics for the curvature perturbation, including the nonlinear relation between $\zeta$ and $\delta\phi$ from our lattice simulations. This would allow us to assess, for example, whether positive non-Gaussianities in $\zeta$ are obtained in all scenarios considered here, as suggested in Firouzjahi and Riotto (2023). Other groups attempted to study the effect of backreaction and nonlinearities in the USR scenarios employing full in-in computations, $\delta N$, or the stochastic inflation formalism (see, e.g., Pattison et al. (2017); Biagetti et al. (2018); Ezquiaga et al. (2020); Ballesteros et al. (2020); Pattison et al. (2021); Cai et al. (2022); Biagetti et al. (2021); Figueroa et al. (2022); Raatikainen et al. (2024); Li et al. (2024); Mizuguchi et al. (2024); Ballesteros and Egea (2024); Animali and Vennin (2024); Jackson et al. (2024)), and it would be interesting to compare these techniques with lattice results. Additionally, evaluating the four-point function of the field fluctuations $\delta\phi$ would provide key insights into the non-perturbative source of SIGW (complementing recent efforts in this direction Cai et al. (2019); Unal (2019); Yuan and Huang (2021); Atal and Domènech (2021); Adshead et al. (2021); Abe et al. (2023); Chang et al. (2023); Garcia-Saenz et al. (2023); Li et al. (2023); Perna et al. (2024); Ruiz and Rey (2024)). These next steps will allow us to better quantify the importance of nonlinearities in the USR scenarios, paving the way for robust observational predictions for GW experiments, as well as constraints on the PBH abundance.